# Incongruent Melting and Phase Diagram of SiC from Machine Learning Molecular Dynamics

### v1

@@ Abstract
Silicon carbide (SiC) is an important technological material, but its high-temperature phase diagram has remained unclear due to conflicting experimental results about congruent versus incongruent melting. Here, we employ large-scale machine learning molecular dynamics (MLMD) simulations to gain insights into SiC decomposition and phase transitions. Our approach relies on a Bayesian active learning workflow to efficiently train an accurate machine learning force field on density functional theory data. Our large-scale simulations provide direct evidence of the incongruent melting of SiC, confirming its decomposition into silicon-rich and carbon phases at high temperature and pressure. During cooling at high pressures, carbon nanoclusters nucleate and grow within the homogeneous molten liquid. During heating, the decomposed mixture reversibly transitions back into a homogeneous SiC liquid. The full pressure-temperature phase diagram of SiC is systematically constructed using MLMD simulations, providing new understanding of the nature of phases, resolving long-standing inconsistencies from previous experiments and yielding technologically relevant implications for processing and deposition of this material.
@@ 1 Introduction
Silicon carbide (SiC) is an important technological material valued for its high hardness, mechanical strength, high thermal conductivity, and wide band gap [1]. The high-temperature and high-pressure behavior of SiC is of great scientific interest for understanding planetary interiors and stellar processes, due to its identification from adsorption spectroscopy of carbon-rich extrasolar planets [2, 3]. Understanding the properties of SiC under extreme conditions is also crucial for various applications, including nuclear reactors [4], epitaxial deposition growth of SiC [5, 6], and graphene synthesis [7] for electronics and quantum devices.
@@
The melting and decomposition of SiC have been investigated by experimental and computational methods for decades. However, various experiments have reached inconsistent conclusions and thereby raised confusion about the nature of melting of SiC. Experiments at high temperatures and pressures are challenging because it is difficult to directly observe the kinetics of melting, and factors such as sample size and purity, and thermal gradients causing temperature inhomogeneity can further complicate interpretations. In some studies, incongruent melting has been reported, with the observation that SiC decomposes into a silicon rich liquid and solid carbon upon heating [8, 9, 10, 11]. Most of these experiments are performed at pressures below 10 GPa, except for a more recent study with a laser-heated diamond-anvil cell that reaches up to 81 GPa and 3200 K [10]. However, other experimental studies report that SiC melts congruently, without decomposition into different phases [12, 13].
@@
Molecular dynamics (MD) simulations using density functional theory (ab initio MD, or AIMD) and various empirical potentials were employed to investigate the melting process of SiC with computational simulations. Phase separation and decomposition were not observed, possibly due to the limitation of the small supercell in AIMD [14, 15]. The AIMD results showed some carbon aggregations within amorphous SiC, resulting in a carbon-carbon (C-C) peak at a graphite / diamond bond length of 1.5 Å in their radial distribution functions (RDF). However, these simulations mainly reveal a dominance of Si-C bonds in the system, with only a minor presence of nanoclusters, and do not capture any decomposition according to the RDF. On the other hand, empirical potentials, such as the Vashishta potential [16], the Tersoff potential [17, 18, 19], and the Gao-Weber potential [20], provide differing descriptions of uncertain accuracy. For example, simulations using the Vashishta potential show no formation of C-C bonds in either liquid or amorphous SiC. Similar to AIMD, the Tersoff potentials [17, 18], the Gao-Weber potential [20], and a recent Behler-Parrinello neural network potential [21] studied the melting of SiC at high temperature and low to moderate pressure (0-10 GPa), showing RDF characteristics where Si-C bonds dominate, with minor C-C bonds signatures. Therefore, the decomposition and phase separation in the amorphous phase of SiC and its melting process remain underexplored and inconsistent across ab initio, empirical, and machine learning atomistic simulations, as well as experimental observations.
@@
To overcome the limitations of previous studies, we employ a machine learning force field (MLFF) trained on density functional theory (DFT) data. First, we collect DFT data for different SiC phases and train an MLFF to describe atomic interactions using a hierarchical Bayesian active learning workflow. This approach efficiently explores the phase space across different temperatures, pressures, and compositions. Then, we perform large-scale molecular dynamics simulations, enabled by GPU acceleration, to reliably model the melting and decomposition of SiC at a sufficient scale. By conducting two-phase coexistence simulations, where the interface between two phases of interest is created and equilibrated, we quantitatively determine, for the first time, the transition temperatures between phases at different pressures.
@@
Our simulations conclusively demonstrate the occurrence of incongruent melting at high pressure and the decomposition of SiC into Si and C phases. The transition temperatures between the zinc-blende SiC, decomposed incongruent mixture, and the homogeneous liquid phase are identified by our simulations, at pressures ranging from 30 GPa to 120 GPa. We obtain the pressure-temperature phase diagram depicting the stability regions of different phases. Notably, our findings of the decomposition indicate that amorphous SiC can only be produced through irradiation, not via melt-quench processes, consistent with existing experimental approaches [22, 23]. Our atomic-level results provide microscopic insights into the decomposition and melting behavior of this technologically important material, resolving discrepancies in previous experimental and computational studies.
2 Methods
Machine learning force field and Bayesian active learning.
Accurately modeling the melting and decomposition of SiC at high temperature and pressure requires an interatomic potential that captures complex bonding environments, including phase separation. Traditional empirical force fields, such as Tersoff [18] or Vashishta [16] potentials, often fail to accurately describe bond breaking and phase transitions. While AIMD [14, 15] provides accurate atomic forces, its high computational cost limits simulations to small system sizes and short timescales. To overcome these challenges, we employ a MLFF trained on DFT calculations, such that the atomic interactions can be modeled with high accuracy, while remaining cost effective for large-scale simulations.
Our MLFF is built using the FLARE framework [24, 25], which is based on a sparse Gaussian process model [26] with atomic cluster expansion descriptors[27]. This model predicts forces while providing uncertainty quantification. To efficiently generate training data, we employ Bayesian active learning, which systematically selects new atomic configurations where the MLFF is most uncertain. The workflow follows these steps: (1) Initial Training: The MLFF is first trained on a small set of DFT-calculated structures, including perfect and slightly perturbed crystal structures. (2) Uncertainty-Guided MD: The MLFF runs MD simulations, predicting atomic forces while simultaneously estimating its uncertainty. (3) Selective DFT Calculations: If the model’s uncertainty exceeds a threshold, new configurations are labeled using DFT. Otherwise, continue with MD. (4) Iterative Refinement: The MLFF is retrained with the newly acquired DFT data, progressively improving its accuracy. This adaptive learning process ensures that the force field is only trained where necessary, avoiding redundant DFT calculations while capturing the relevant phase space. Once the active learning is complete, we map the SGP into an equivalent but more computationally efficient polynomial model [26, 25], which serves as the final MLFF for large-scale production simulations in LAMMPS [28]. The DFT calculations are performed using VASP [29] with the PBE exchange correlation functional [30], following the computational settings in our previous works [31, 25].
The training set is initialized with structures from our previous study [25], encompassing the hexagonal 2H/4H/6H polytypes, B3 (zinc blende or 3C) and the high pressure B1 (rock salt) phases. We then employ the FLARE Bayesian active learning workflow to systematically sample data that captures melting and decomposition behavior across different pressures. Active learning simulations are performed at 0, 30, 60, 90, and 120 GPa over a temperature range of 2000-6000 K until melting is observed. Additional data is collected for pure carbon (diamond and graphite phases, 2000-4000 K) and pure silicon (1000-3000 K) across all pressure conditions. The trained MLFF is then used to investigate SiC melting and decomposition under high temperature and pressure conditions.
Large-scale molecular dynamics simulation.
The large-scale molecular dynamics simulations are conducted using GPU-accelerated LAMMPS with Kokkos parallelization [32] and the FLARE MLFF pairstyle, which has demonstrated exceptional performance in billion-atom catalytic simulations [33]. During the melting simulations, the NPT ensemble with anisotropic (aniso) control of cell dimensions is employed, allowing independent scaling along the $x$, $y$ and $z$ directions through stress components $P_{xx}$, $P_{yy}$, and $P_{zz}$. Phase separation exhibits a tendency to minimize the interface between decomposed phases, necessitating anisotropic elongation of the supercell, as illustrated in Fig.1 and Fig.2.
@@ 3 Results
3.1 Higher-temperature liquid decomposition and phase separation
As described in the Method section, the MLFF was iteratively refined via Bayesian active learning to ensure accuracy across the phase space, while allowing us to simulate large systems efficiently. We performed Bayesian active learning hierarchically [34], starting with a 64-atom supercell, and then extends to a 512-atom supercell. The 64 atom cell is insufficient for capturing the decomposition, while the larger 512-atom supercell enables observation of the phase separation process. As shown in Fig.1, the phase decomposition occurs at the pressure of 30 GPa during the active learning process, prior to which the carbon and silicon atoms form clusters instead of mixing homogeneously. As a result, the active learning simulation collects DFT data from both Si-rich and C-rich regimes. This observation also provides evidence for the existence of incongruent melting and decomposition even in small-scale simulations. Nevertheless, to avoid finite size effects, much larger simulations are necessary to quantify the phase transition temperature.
Figure 1: DFT data collected during active learning simulation with 512 atoms at P = 30 GPa, T = 5000 K.
@@ 3.1 Higher-temperature liquid decomposition and phase separation
Starting with the well-trained MLFF model, we run large-scale heating and cooling simulations at 30, 60, and 80 GPa to investigate the phase transformation and decomposition of SiC. We start with a 512,000-atom bulk crystal supercell of the cubic zinc-blende (ZB/3C/B3) phase. At these pressures, decomposition does not occur readily during the MD simulations due to hysteresis, even when heating the cubic ZB crystal structure to temperatures as high as 4000 K. The crystal structure is only destroyed when the temperature is increased to 5000 K.
@@
Figure 2: Left: decomposed configuration of 512,000 atoms at the end of the simulation at 3000 K and 60 GPa. Yellow: Si, black: C. Middle: atoms colored with lattice type classified by polyhedral template matching. Blue: cubic diamond, orange: hexagonal diamond, purple: graphite, white: others. Right: radial distribution function of the structure.
At ambient pressure (close to 0 GPa), we observe sublimation as the unit cell volume expands, causing the SiC crystal to break into small molecules, consistent with experimental observations. In contrast, at pressures of 30, 60, and 80 GPa, sublimation is not observed.
@@
At 5000 K, small separate carbon and silicon clusters form in the amorphous configurations, which indicates a trend away from a homogeneous mixture of silicon and carbon. Even though the temperature is significantly higher than experimentally reported incongruent or congruent melting points [10, 11, 35, 8, 9], the complete phase separation of C and Si does not occur. Upon cooling the system to 3000 K, we observe decomposition into carbon crystals and silicon liquid. Specifically, as the temperature drops below 4000 K, carbon clusters start to grow in size, while the Si concentration in the liquid phase increases. Upon further cooling to 3000 K, larger C clusters consisting of recognizable graphite or diamond crystal structure form within the Si-rich liquid.
@@
The formation of either graphite or diamond C phases depends on the simulation pressure. The C clusters are classified as graphite-like or diamond-like structures by the polyhedral template matching [36] implemented in OVITO [37]. At 30 GPa, C clusters form graphite-like honeycomb lattice structures, while the Si-rich liquid persists throughout our simulation temperature range (3000-5000 K), which is higher than the Si melting point (1684 K at 0 GPa [38]). At 60 GPa, the diamond and graphite-like clusters incorporate Si atoms within them, as illustrated in Fig.2. In addition, the C-C pair distribution function exhibits a dominant peak around 1.5 Å, while the C-Si peak decreases. At 80 GPa, the decomposition results in the separation of diamond C, liquid Si, and the high-pressure rock-salt (RS/B1) phase of SiC. The configurations at 30 GPa and 80 GPa are shown in the Supplementary Materials. Our results agree with a recent experimental observation [10] that solid C and liquid Si form upon melting (i.e., incongruent melting) of SiC at high pressures. At 120 GPa, the zinc blende to rock salt phase transition occurs, and the high pressure stabilizes the crystal structure and prevents it from amorphization, with no liquid phase appearing up to 5000 K.
@@
3.2 High temperature transition from decomposed Si + C to homogeneous liquid
Having established that decomposition occurs, we next explore the phase transition between the decomposed and homogeneous liquid states. To verify the reversibility of the incongruent high-pressure melting and to quantitatively identify the transition temperature and pressure of the spontaneous decomposition, we set up a smaller MD simulation with slower cooling and heating rates.
@@ 3.2 High temperature transition from decomposed Si + C to homogeneous liquid
Specifically, we start with 8,000 atoms in the zinc-blende (B3) phase of SiC at 5000 K, so that the crystal becomes a homogeneous liquid. With a time step of 0.5 fs, we perform an equilibration for 0.5 ns at each temperature, followed by a 0.5 ns cooling process at a rate of 400 K/ns. We alternate between equilibration and cooling steps until reaching 3000 K, then we ramp the temperature back up to 5000 K following the same scheme. The temperature profile over simulation time is shown as the blue curve in the lower panel of Fig.3.
@@
Figure 3: Cooling and heating simulation of SiC at 30 GPa, illustrating the transition between decomposed Si + C and homogeneous liquid. Top: population distribution of local C atom fraction with the simulation time. Red: high population, blue: low population. Bottom: temperature (blue) and largest carbon cluster size (red) (number of atoms) with the simulation time. Blue shadows indicate the phase transition regions.
@@
To quantitatively identify spontaneous phase separation, we use the local C fraction (concentration) as the order parameter and track the size of the largest C cluster. At each time step, we divide the simulation cell into $10\times 10\times 10$ voxels and calculate the C atom fraction in each voxel. The population of voxels with different local fraction levels over time is represented by the color scale in the top panel of Fig.3, with red indicating higher populations.
@@
At the beginning of the simulation, at a pressure under 30 GPa and a high temperature of 5000 K, we observe a high population of 0.5 local C fraction values, indicating a homogeneous liquid with equal mixture of Si and C atoms. The lower panel of Fig.3 shows the system temperature and the largest C cluster size. As the system cools, the largest C cluster size increases, notably around 4000 K. At 7 ns, when the system cools down to 3600 K, spontaneous decomposition is finished, with C atoms mostly depleted from the Si liquid. The local C-fractions are either near 0 (Si rich) or near 1 (C rich), as indicated in the upper panel.
@@
During heating, the reverse process initiates at the same transition temperature of approximately 4000 K. The C cluster dissolves into the Si liquid, and the local C fraction returns to around 0.5, analogous to the initial state of the simulation. This indicates that the system has returned to the homogeneous liquid upon heating. Similar results are obtained for 60 GPa and 90 GPa, which are shown in the Supplementary Materials.
3.3 Lower temperature transition from crystalline to decomposed Si + C
Having identified the high-temperature transition between the incongruently decomposed Si + C mixture and the homogeneous liquid phase, we investigate the lower-temperature transition between the crystal B1 or B3 phases and the decomposed phase. To overcome hysteresis and high nucleation barriers that prevent decomposition in direct MD simulations of B3 SiC crystals heated to 4000 K, we employ two-phase coexistence simulations. This approach, comprising four stages [39, 40], allows us to quantitatively determine the transition temperatures: 1) Equilibrate the crystalline supercell at pressure $P$ (30, 60, 90 GPa) and an initially estimated transition temperature $T_{es}$. 2) Fix the positions and velocities of half the atoms in a slab along z-axis and heat the other half up to $T_{h}$ (much higher than $T_{es}$) in the $NP_{zz}T$ ensemble to induce a transition to the liquid phase. 3) Cool the heated region from $T_{h}$ to $T_{es}$ in the $NP_{zz}T$ ensemble, forming the decomposed phase while maintaining the unheated crystalline half. 4) Remove constraints on the crystalline region and equilibrate the entire system in the $NPH$ ensemble, allowing the phase boundary to move. In stages 1-3, we create an interface between the crystalline and decomposed phases. In the last stage, the system equilibrates near the transition temperature. If both phases coexist at equilibrium, we identify this temperature as the phase transition temperature.

### v2

@@ Abstract
Silicon carbide (SiC) is an important technological material, but its high-temperature phase diagram has remained unclear due to conflicting experimental results about congruent versus incongruent melting. Here, we employ large-scale machine learning molecular dynamics (MLMD) simulations to gain insights into SiC decomposition and phase transitions. Our approach relies on a Bayesian active learning workflow to efficiently train an accurate machine learning force field on density functional theory data. Our large-scale simulations provide direct indication that melting of SiC proceeds incongruently via decomposition into silicon-rich and carbon phases at high temperature and pressure. During cooling at high pressures, carbon nanoclusters nucleate and grow within the homogeneous molten liquid. During heating, the decomposed mixture reversibly transitions back into a homogeneous SiC liquid. The full pressure-temperature phase diagram of SiC is systematically constructed using MLMD simulations, providing new understanding of the nature of phases, resolving long-standing inconsistencies from previous experiments and yielding technologically relevant implications for processing and deposition of this material.
@@ 1 Introduction
Silicon carbide (SiC) is an important technological material valued for its high hardness, mechanical strength, high thermal conductivity, and wide band gap [1]. The high-temperature and high-pressure behavior of SiC is of great scientific interest for understanding planetary interiors and stellar processes, due to its identification from adsorption spectroscopy of carbon-rich extrasolar planets [2]. Understanding the properties of SiC under extreme conditions is also crucial for various applications, including nuclear reactors [3], epitaxial deposition growth of SiC [4, 5], and graphene synthesis [6] for electronics and quantum devices.
@@
The melting and decomposition of SiC have been investigated by experimental and computational methods for decades. However, various experiments have reached inconsistent conclusions and thereby raised confusion about the nature of melting of SiC. Experiments at high temperatures and pressures are challenging because it is difficult to directly observe the kinetics of melting, and factors such as sample size and purity, and thermal gradients causing temperature inhomogeneity can further complicate interpretations. Moreover, internal inconsistencies within individual studies and substantial disagreement between independent measurements indicate that the experimental picture remains under-established.
In some studies, incongruent melting has been reported, with the observation that SiC decomposes into a silicon rich liquid and solid carbon upon heating [7, 8, 9, 10]. Reported incongruent melting (decomposition) onset temperatures span a broad range: multiple sets of measurements near 10 GPa clusters in the 2800–3500 K range [10, 11, 12, 8, 7], whereas a high-pressure, laser-heated diamond-anvil cell (DAC) study [9] reports substantially lower decomposition temperatures ($\sim 2000$ K), creating a discontinuity with the lower-pressure results. This dispersion, compounded by conflicting reports of congruent melting [13, 14], highlights that the SiC decomposition phase boundary remains experimentally under-determined and motivates a qualitative and quantitative re-examination.
@@
Molecular dynamics (MD) simulations using density functional theory (ab initio MD, or AIMD) and various empirical potentials were employed to investigate the melting process of SiC with computational simulations. AIMD studies [15, 16] do not reach extended Si-rich / C-rich phase separation, although at the highest reported temperature (11000 K) the C–C peak in the radial distribution function (RDF) begins to overtake the Si–C peak, indicating substantial carbon clustering and reduction of silicon-carbon mixing. This captures the qualitative tendency toward decomposition, but the need for unrealistically high temperature together with limited cell size and timescale prevents quantitative determination of the true decomposition boundary or formation of fully phase-separated domains. On the other hand, empirical potentials, such as the Vashishta potential [17], the Tersoff potential [18, 19, 20], and the Gao-Weber potential [21], provide differing descriptions of uncertain accuracy. For example, simulations using the Vashishta potential show no formation of C-C bonds in either liquid or amorphous SiC (see Supplementary Section 6.1). Similarly, the Tersoff potentials [18, 19], the Gao-Weber potential [21], and a recent Behler-Parrinello neural network potential [22] studied the melting of SiC at high temperature and low to moderate pressure (0–10 GPa), showing RDF characteristics where Si-C bonds dominate, with minor C-C bonds signatures. More recent SiC ML interatomic potentials target relevant phenomena, including a UF3 model for high-temperature surface sublimation at near-ambient pressures [23] and a deep learning potential optimized for irradiation damage cascades [24]. But they are not designed or validated for the extended high-pressure (>10>10 GPa) solid–liquid decomposition transformations. Therefore, the decomposition and phase separation in the amorphous phase of SiC and its melting process remain underexplored and inconsistent across ab initio, empirical, and machine learning atomistic simulations, as well as experimental observations.
@@
To overcome the limitations of previous studies, we employ a machine learning force field (MLFF) trained on density functional theory (DFT) data. First, we collect DFT data for different SiC phases and train an MLFF to describe atomic interactions using a hierarchical Bayesian active learning workflow. This approach efficiently explores the phase space across different temperatures, pressures, and compositions. Then, we perform large-scale molecular dynamics simulations, enabled by GPU acceleration, to reliably model the melting and decomposition of SiC at a sufficient scale. By conducting two-phase coexistence simulations, where the interface between two phases of interest is created and equilibrated, we quantitatively determine, for the first time, the temperatures and pressures of transitions between phases.
@@
Our simulations provide strong evidence for incongruent melting at high pressure and offer insights into the mechanism of SiC decomposition into Si and C rich phases. The transition temperatures between the zinc-blende SiC, decomposed incongruent mixture, and the homogeneous liquid phase are identified by our simulations, at pressures ranging from 10 GPa to 120 GPa. We obtain the pressure-temperature phase diagram depicting the stability regions of different phases. Notably, our findings of the decomposition indicate that amorphous SiC can only be produced through irradiation, and not via melt-quench processes [25]. Our atomic-level results provide microscopic insights into the decomposition and melting behavior of this technologically important material, resolving discrepancies in previous experimental and computational studies.
@@ 2 Results
2.1 Bayesian active learning of machine learning force field
Our MLFF is based on Gaussian process regression, which provides both force predictions and per-atom uncertainty estimates [26, 27]. This enables a Bayesian active learning workflow (Fig. 1a) where MD simulations are propagated with the current surrogate, and configurations whose uncertainties exceed a threshold trigger single-point DFT queries for training data collection and surrogate update. As more details described in the Methods section, the MLFF was iteratively refined via Bayesian active learning to ensure accuracy across the phase space, while allowing us to simulate large systems efficiently. We performed multiple Bayesian active learning trajectories in parallel for different compositions (pure Si, pure C, and SiC) at various temperatures and pressures (Fig. 1b).
While a 64-atom cell is insufficient to capture decomposition, the larger 512-atom supercell enables observation of the phase separation process. As shown in Fig. 1d, the first Si-rich and C-rich domains emerged spontaneously at 30 GPa during unbiased high-$T$ / high-$P$ active learning MD, prior to any decomposed structures being present in the training set. We did not seed or impose this Si and C phase separation. Instead, trajectories were advanced with the current surrogate, and single-point DFT labels were added only when Bayesian force uncertainties on naturally visited clustered configurations exceeded the acquisition threshold. This procedure preserves high-fidelity force predictions on-the-fly and expands the training set into newly accessed regions of configuration space. Consequently, the database expanded to include Si-rich and C-rich regimes solely through the natural evolution of system’s dynamics. This spontaneous appearance provides independent evidence that incongruent melting (decomposition) occurs without being assumed a priori, even in modest cell sizes.
Nevertheless, the phase separation observed in the 512-atom active learning simulations does not provide comprehensive information about the decomposition process. For instance, the limited cell size is insufficient to reveal whether the C and Si clusters form crystalline or liquid phases, or to determine the precise conditions under which decomposition occurs. To qualitatively characterize the nature of these phases and their structural properties, we perform large-scale MD simulations with up to 512,000 atoms (Sec. 2.2). Furthermore, to quantitatively determine the precise phase transition temperatures and pressures, we employ two-phase coexistence simulations (Sec. 2.3 and Sec. 2.4).
Figure 1: a. Bayesian active learning workflow for collecting DFT training data on the fly during MD simulations. The MLFF predicts forces and uncertainties; configurations exceeding the uncertainty threshold trigger DFT labeling. b. Training data composition and coverage. Multiple Bayesian active learning trajectories were performed in parallel for different compositions (pure Si, pure C, and SiC) across various temperatures and pressures. For SiC, we started with a 64-atom supercell and then extended to a 512-atom supercell where phase separation spontaneouly occurred. c. Two-phase simulation protocol to determine phase transition temperatures. Interfaces between the solid/liquid and decomposed phases were created by inducing decomposition, followed by a two-phase coexistence simulation in the NPH ensemble to converge on the final transition temperature. d. Spontaneous decomposition during active learning. The collected data includes states where SiC decomposition occurred spontaneously during the active learning MD simulations, prior to any decomposed structures being present in the training set.
2.2 Incongruent melting from large-scale MD
@@
Starting with the well-trained MLFF model, we run large-scale heating and cooling simulations at 30, 60, and 80 GPa to qualitatively characterize the nature of the phase transformation and decomposition of SiC. We start with a 512,000-atom bulk crystal supercell of the cubic zinc-blende (ZB/3C/B3) phase. At these pressures, decomposition does not occur readily during the MD simulations due to hysteresis, even when heating the cubic ZB crystal structure to temperatures as high as 4000 K. The crystal structure is only destroyed when the temperature is increased to 5000 K.
@@
Figure 2: Left: decomposed configuration of 512,000 atoms at the end of the simulation at 3000 K and 60 GPa. Yellow: Si, black: C. Middle: atoms colored with lattice type classified by polyhedral template matching. Blue: cubic diamond, orange: hexagonal diamond, purple: graphite, white: others. Right: radial distribution function of the structure shows the C-C peak dominates over the Si-C peak.
@@
At 5000 K across 30, 60 and 80 GPa, small separate carbon and silicon clusters form in the amorphous configurations, which indicates a trend away from a homogeneous mixture of silicon and carbon. Even though the temperature is significantly higher than experimentally reported incongruent or congruent melting points [9, 10, 11, 7, 8], the complete phase separation of C and Si does not occur. Upon cooling the system to 3000 K, we observe decomposition into carbon crystals and silicon liquid. Specifically, as the temperature drops below 4000 K, carbon clusters start to grow in size, while the Si concentration in the liquid phase increases. Upon further cooling to 3000 K, larger C clusters consisting of recognizable graphite or diamond crystal structure form within the Si-rich liquid.
@@
The formation of either graphite or diamond C phases depends on the simulation pressure. The C clusters are classified as graphite-like or diamond-like structures by the polyhedral template matching [28] implemented in OVITO [29]. At 30 GPa, C clusters form graphite-like honeycomb lattice structures, while the Si-rich liquid persists throughout our simulation temperature range (3000–5000 K), which is higher than the Si melting point (1684 K at 0 GPa [30]). At 60 GPa, the diamond and graphite-like clusters incorporate Si atoms within them, as illustrated in Fig. 2. In addition, the C-C pair distribution function exhibits a dominant peak around 1.5 Å, while the C-Si peak decreases. At 80 GPa, the decomposition results in the separation of diamond C, liquid Si, and the high-pressure rock-salt (RS/B1) phase of SiC. We have included the configurations at 30 GPa and 80 GPa and more structural analysis details in Supplementary Section 6.2 and 6.3.
Our results agree with a recent experimental observation [9] that solid C and liquid Si form upon melting (i.e., incongruent melting) of SiC at high pressures. At 120 GPa, the zinc blende to rock salt phase transition occurs, and the high pressure stabilizes the crystal structure and prevents it from amorphization, with no liquid phase appearing up to 5000 K.
In the next sections, we quantitatively determine the transition temperatures between the crystalline, decomposed, and homogeneous liquid phases using two-phase coexistence simulations, and build the P-T phase diagram of SiC.
@@
2.3 High-temperature phase boundary: decomposed Si+C $\leftrightarrow$ homogeneous liquid
@@
Having established that decomposition occurs, we next explore the phase transition between the decomposed and homogeneous liquid states. To verify the reversibility of the incongruent high-pressure melting and to quantitatively identify the transition temperature and pressure of the spontaneous decomposition, we set up a smaller MD simulation with slower cooling and heating rates. Specifically, we start with 8,000 atoms in the zinc-blende (B3) phase of SiC at 5000 K, so that the crystal becomes a homogeneous liquid. A conservative simulation time step of 0.5 fs was selected due to the elevated temperatures and pressures, the short vibrational periods of light Si and C atoms, and the need to accurately capture rapid bond rearrangements during decomposition while minimizing integration error. We perform an equilibration for 0.5 ns at each temperature, followed by a 0.5 ns cooling process at a rate of 400 K/ns. We alternate between equilibration and cooling steps until reaching 3000 K, then we ramp the temperature back up to 5000 K following the same scheme. The temperature profile over simulation time is shown as the blue curve in the middle panel of Fig. 3.
@@
Figure 3: SiC phase transitions and coexistence at 30 GPa from MD simulations. Top and middle panels (from 8,000-atom cooling and heating simulations) illustrate the reversibility of phase transition. Top: population distribution of local C atom fraction over time: a single peak indicates the homogeneous liquid phase, while multiple peaks indicate the decomposed (Si+C) phase (red: high population, blue: low population). Middle: temperature (blue) and largest carbon cluster size (red, number of atoms) with the simulation time, with blue shaded regions indicating the phase transition boundaries. Bottom panel (from 16,000-atom two-phase simulations) confirms the coexistence temperature by tracking the largest carbon cluster size (left) and temperature convergence at the NPHNPH stage (right).
@@
To identify spontaneous phase separation, we use the local C fraction (concentration) as the order parameter and track the size of the largest C cluster. At each time step, we divide the simulation cell into $10\times 10\times 10$ voxels and calculate the C atom fraction in each voxel. The population of voxels with different local fraction levels over time is represented by the color scale in the top panel of Fig. 3, with red indicating higher populations.
@@
At the beginning of the simulation, at a pressure under 30 GPa and a high temperature of 5000 K, we observe a high population of 0.5 local C fraction values, indicating a homogeneous liquid with equal mixture of Si and C atoms. The middle panel of Fig. 3 shows the system temperature and the largest C cluster size. As the system cools, the largest C cluster size increases, notably around 4000 K. At 7 ns, when the system cools down to 3600 K, spontaneous decomposition is finished, with C atoms mostly depleted from the Si liquid. The local C-fractions are either near 0 (Si rich) or near 1 (C rich), as indicated in the upper panel.
@@
During heating, the reverse process initiates at the same transition temperature of approximately 4000 K. The C cluster dissolves into the Si liquid, and the local C fraction returns to around 0.5, analogous to the initial state of the simulation. This indicates that the system has returned to the homogeneous liquid upon heating. Similar results are obtained for 60 GPa and 90 GPa, which are shown in the Supplementary Section 7.1.
Having confirmed the reversibility, we proceed with a quantitative estimation of the transition temperature between the homogeneous liquid and decomposed Si + C phase, with a four-stage simulation protocols shown in Fig. 1c:
  1. 1.
    Initial Equilibration: The crystalline supercell is first melted and then cooled to the estimated transition temperature $T_{es}$. The entire cell is then equilibrated at this temperature to prepare the initial liquid phase.
  2. 2.
    Interface Creation: To create an interface, the positions and velocities of half the atoms in a slab along $z$-axis are fixed, while the other half is cooled further to a lower temperature, $T_{l}$ (choosen here as 3500 K) to induce the formation of the decomposed phase.
  3. 3.
    Interface Refinement: The cooled region is heated back from $T_{l}$ to $T_{es}$ in the $NP_{zz}T$ ensemble. This action ensures a minimal thermal gradient across the cell while creating a stable interface between the decomposed phase and the liquid phase.
  4. 4.
    NPHNPH Coexistence: Constraints are removed, and the entire system is equilibrated in the $NPH$ ensemble. The phase boundary equilibrates, and the system temperature converges to the phase transition temperature.
The first three stages prepare the two-phase interface near the estimated transition point. During the final $NPH$ stage, the system equilibrates. The temperature at which both phases coexist in equilibrium is then determined as the phase transition temperature.
As shown in bottom panels of Fig. 3, the initial liquid phase at $T_{es}$ is prepared after 0.5 ns (0.2 ns melting, 0.1 ns cooling, and 0.2 ns equilibration). Then, the decomposed phase is clearly formed during stage 2 and 3 (achieved by 0.25 ns of cooling followed by 0.25 ns of heating), indicated by the sharp increase in the size of the largest carbon cluster. Starting at the 1 ns mark, the system enters the final $NPH$ coexistence stage and subsequently converges to a final transition temperature of 4164 K in the coexistence simulation. We note that for runs initialized with an estimated temperature of $T_{es}=4300K$, the temperature was too high to stablize the decomposed phase, and temperature convergence was therefore not achieved. Thus, we computed the final average transition temperature from the remaining three converged runs. Additional two-phase simulation results for pressures at 45, 60, 70 GPa are provided in the Supplementary Section 7.1.
2.4 Low-temperature phase boundary: crystal $\leftrightarrow$ decomposed Si+C
Next, we investigate the lower-temperature transition between the crystal (B1 or B3) phases and the decomposed phase. To overcome hysteresis and high nucleation barriers that prevent decomposition in direct MD simulations of B3 SiC crystals heated to 4000 K, we employ a similar four-stage two-phase coexistence simulation approach [31], as illustrated in Fig. 1c.
  1. 1.
    Initial Equilibration: Equilibrate the crystalline supercell at pressure $P$ (10-90 GPa) and an initially estimated transition temperature $T_{es}$.
  2. 2.
    Interface Creation: Fix half of the $z$-slab, while melt the other half by heating it to high temperature $T_{h}\gg T_{es}$ in the $NP_{zz}T$ ensemble to induce the liquid phase.
  3. 3.
    Interface Refinement: The melted half is then cooled from $T_{h}$ to $T_{es}$ in the $NP_{zz}T$ ensemble. This process facilitates the formation of the decomposed phase while maintaining the unperturbed crystalline half, thus creating a stable crystal-decomposed interface.
  4. 4.
    NPHNPH Coexistence: Constraints are removed, and the entire system is equilibrated in the $NPH$ ensemble. The phase boundary equilibrates, and the system temperature converges to the phase transition temperature.